# Distributed Hypothesis Testing Under A Covertness Constraint

### v1

@@ I Introduction
In this work, the sensor again communicates to the decision center over $n$ uses of a DMC. However, under $\mathcal{H}=0$, communication needs to be such that an external warden cannot detect the mere fact that communication is happening. Depending on the practical situation, it might be important that communication be covert under both hypothesie, or only in the case where there is no alert situation, i.e., under $\mathcal{H}=0$. In this paper, we consider the latter scenario. Implicitly, such a covertness constraint imposes that under $\mathcal{H}=0$, the sensor employs most of the time a dedicated zero-symbol that also represents the absence of communication. It can be shown that the covertness constraint is weaker than imposing that the channel can only be used a sublinear number of times but stronger than the expected cost constraint. However, as is common for covert communication, here we allow the sensor and the decision center to share a uniform secret key that is unknown to the warden, see Figure 1.
@@
where $n_{x^{n},y^{n}}(a,b)$ is the number of occurrences of the pair $(a,b)$ in $(x^{n},y^{n})$. We use $\mathcal{T}_{\mu}^{(n)}(P_{XY})$ to denote the jointly strongly typical set as defined in [21, Definition 2.9].
@@ III Results
$E_{2}$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})\end{subarray}}D\!\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13b)
$E_{3}$ $\triangleq$ $\max_{\begin{subarray}{c}x_{1}\in\mathcal{X}\setminus\{0\}\end{subarray}}\Big[\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\in\bar{\mathcal{P}}_{U}\end{subarray}(x_{1})}D\!\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13c)
$\hskip 85.35826pt+\ D(\Gamma_{Y|X=0}\|\Gamma_{Y|X=x_{1}})\Big],$
@@
    $\theta_{\textnormal{covert}}^{*}(\epsilon)$ $\geq$ $\min\big\{E_{2},E_{3}\big\}.$ (17)
@@
We observe that if $T_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})$ for all $x_{1}$, then $E_{2}=D(P_{V}\|Q_{V})$, which coincides with the local exponent at the decision center without communication from the sensor.
In particular, $E_{2}=D(P_{V}\|Q_{V})$ when $T_{U}=P_{U}$ (as is the case for testing against independence where $Q_{UV}=P_{U}P_{V}$). For $T_{U}=P_{U}$ also $E_{1}=D(P_{V}\|Q_{V})$ and thus communication from the sensor cannot improve the Stein-exponent even for DMCs satisfying (15).
@@
The second summand is strictly positive, except when $\pi_{U|V}(u|v)=Q_{U|V}(u|v)$ for all $u$ and all $v$ with $P_{V}(v)>0$. Thus, to deduce that $E_{1}>D(P_{V}\|Q_{V})$ or $E_{2}>D(P_{V}\|Q_{V})$ it suffices to prove that $\pi_{U|V}=Q_{U|V}$ is not a valid choice in the minimizations. Inspecting the two minimizations, we see that $\pi_{U|V}=Q_{U|V}$ is not a permissible choice in the minimization of $E_{1}$ when $T_{U}\neq P_{U}$ and of $E_{2}$ when $T_{U}\in\ \bar{\mathcal{P}}_{U}(x_{1})$ for some $x_{1}$.
Notice also that $E_{3}>D(P_{V}\|Q_{V})$ by Assumptions (7). ∎
@@
$E_{2}$ $=$ $\min_{\pi_{U}\notin\bar{\mathcal{P}}_{U}(1)}D(\pi_{U}\|Q_{U})=0.2095$ (29)
$E_{3}$ $=$ $D(\Gamma_{Y|X}(\cdot|0)\|\Gamma_{Y|X}(\cdot|1))=0.1170.$ (30)
@@ IV-A Achievability of (16)
where in $(a)$ the second part of Lemma 1 and [21, Remark to Lemma 2.12].
@@ V Converse result in (16)
Notice that by Lemma 2 the cardinality of $|\tilde{\mathcal{X}}^{n}|$ grows subexponentially in $n$ and thus the ratio $\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}$ is of the form $e^{-n\phi_{n}}$ for a vanishing sequence $\phi_{n}$. I.e., for any $\bar{s}\in\mathcal{\bar{S}}_{\bar{x}^{n}}$:
@@
$\phi_{n}:=\log\frac{c|\tilde{\mathcal{X}}^{n}|}{1-\epsilon-\eta}.$ (74)
@@
$=$ $\Pr\left[S\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]Bigg(1-\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\Bigg)+\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|},$ (83)
@@ Appendix C Proof of Lemma 3
By (118) and the blowing-up lemma [22, Remark on p. 446]:
@@
Let $\{\mu_{n}\}_{n=1}^{\infty}$ denote a sequence of positive numbers satisfying [^1] Condition (125b) ensures that the probability of the strongly typical set $\mathcal{T}_{\mu_{n}}^{(n)}(P_{XY})$ under $P_{XY}^{\otimes n}$ converges to one as $n\to\infty$; see [21, Remark following Lemma 2.12].

### v2

@@ I Introduction
In this work, the sensor again communicates to the decision center over $n$ uses of a DMC. However, under $\mathcal{H}=0$, communication needs to be such that an external warden cannot detect the mere fact that communication is happening. Depending on the practical situation, it might be important that communication be covert under both hypotheses, or only in the case where there is no alert situation, i.e., under $\mathcal{H}=0$. In this paper, we consider the latter scenario. This asymmetric requirement is motivated by practical monitoring systems where no transmission should be detected under normal conditions $(\mathcal{H}=0)$, while when an abnormal event occurs $(\mathcal{H}=1)$, transmitting and thus being detected is acceptable. Implicitly, such a covertness constraint imposes that under $\mathcal{H}=0$, the sensor employs most of the time a dedicated zero-symbol that also represents the absence of communication. It can be shown that the covertness constraint is weaker than imposing that the channel can only be used a sublinear number of times but stronger than the expected cost constraint. However, as is common for covert communication, here we allow the sensor and the decision center to share a uniform secret key that is unknown to the warden, see Figure 1.
@@
where $n_{x^{n},y^{n}}(a,b)$ is the number of occurrences of the pair $(a,b)$ in $(x^{n},y^{n})$. We use $\mathcal{T}_{\mu}^{(n)}(P_{XY})$ to denote the jointly strongly typical set as defined in [7, Definition 2.9].
@@ III Results
$E_{2}(x_{1})$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})\end{subarray}}D\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13b)
$E_{3}(x_{1})$ $\triangleq$ $\min_{\begin{subarray}{c}\pi_{UV}:\ \\ \pi_{V}=P_{V},\;\pi_{U}\in\bar{\mathcal{P}}_{U}(x_{1})\end{subarray}}D\big(\pi_{UV}\,\|\,Q_{UV}\big)$ (13c)
$\hskip 85.35826pt+\ D(\Gamma_{Y|X=0}\|\Gamma_{Y|X=x_{1}}),$
@@
    $\theta_{\textnormal{covert}}^{*}(\epsilon)$ $\geq$ $\max_{\begin{subarray}{c}x_{1}\in\mathcal{X}\setminus\{0\}\end{subarray}}\min\big\{E_{2}(x_{1}),E_{3}(x_{1})\big\}.$ (17)
@@
We observe that if $T_{U}\notin\bar{\mathcal{P}}_{U}(x_{1})$ for all $x_{1}$, then $E_{2}(x_{1})=D(P_{V}\|Q_{V})$, which coincides with the local exponent at the decision center without communication from the sensor.
In particular, $E_{2}(x_{1})=D(P_{V}\|Q_{V})$ when $T_{U}=P_{U}$ (as is the case for testing against independence where $Q_{UV}=P_{U}P_{V}$). For $T_{U}=P_{U}$ also $E_{1}=D(P_{V}\|Q_{V})$ and thus communication from the sensor cannot improve the Stein-exponent even for DMCs satisfying (15).
@@
The second summand is strictly positive, except when $\pi_{U|V}(u|v)=Q_{U|V}(u|v)$ for all $u$ and all $v$ with $P_{V}(v)>0$. Thus, to deduce that $E_{1}>D(P_{V}\|Q_{V})$ or $E_{2}(x_{1})>D(P_{V}\|Q_{V})$ it suffices to prove that $\pi_{U|V}=Q_{U|V}$ is not a valid choice in the minimizations. Inspecting the two minimizations, we see that $\pi_{U|V}=Q_{U|V}$ is not a permissible choice in the minimization of $E_{1}$ when $T_{U}\neq P_{U}$ and of $E_{2}(x_{1})$ when $T_{U}\in\ \bar{\mathcal{P}}_{U}(x_{1})$ for some $x_{1}$.
Notice also that $E_{3}(x_{1})>D(P_{V}\|Q_{V})$ by Assumptions (7). ∎
@@
$E_{2}(1)$ $=$ $\min_{\pi_{U}\notin\bar{\mathcal{P}}_{U}(1)}D(\pi_{U}\|Q_{U})=0.2095$ (29)
$E_{3}(1)$ $=$ $D(\Gamma_{Y|X}(\cdot|0)\|\Gamma_{Y|X}(\cdot|1))=0.1170.$ (30)
@@ IV-A Achievability of (16)
where in $(a)$ the second part of Lemma 1 and [7, Remark to Lemma 2.12].
@@ V Converse result in (16)
Notice that by Lemma 2 the cardinality of $|\tilde{\mathcal{X}}^{n}|$ grows sub-exponentially in $n$ and thus the ratio $\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}$ is of the form $e^{-n\phi_{n}}$ for a vanishing sequence $\phi_{n}$. I.e., for any $\bar{s}\in\mathcal{\bar{S}}_{\bar{x}^{n}}$:
@@
$\phi_{n}:=\frac{1}{n}\ln\frac{c|\tilde{\mathcal{X}}^{n}|}{1-\epsilon-\eta}.$ (74)
@@
$=$ $\Pr\left[S\in\bar{\mathcal{S}}_{\bar{x}^{n}}\right]\Bigg(1-\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|}\Bigg)+\frac{1-\epsilon-\eta}{c|\tilde{\mathcal{X}}^{n}|},$ (83)
@@ Appendix C Proof of Lemma 3
By (118) and the blowing-up lemma [12, Remark on p. 446]:
@@
Let $\{\mu_{n}\}_{n=1}^{\infty}$ denote a sequence of positive numbers satisfying [^1] Condition (125b) ensures that the probability of the strongly typical set $\mathcal{T}_{\mu_{n}}^{(n)}(P_{XY})$ under $P_{XY}^{\otimes n}$ converges to one as $n\to\infty$; see [7, Remark following Lemma 2.12].